# LATTICES OF FLATS FOR SYMPLECTIC MATROIDS

### v5

@@ Abstract.
We are interested in expanding our understanding of symplectic matroids by exploring the properties of a class of symplectic matroids with a "lattice of flats". Taking a well-behaved family of subdivisions of the cross polytope we obtain a construction of lattices, resembling a known definition for the geometric lattice corresponding to ordinary matroid. We construct a correspondence to a set of enveloped symplectic matroids, we denote ranked symplectic matroids. As a byproduct of our construction, we also obtain a new way of finding symplectic matroids from ordinary ones and an embedding theorem into geometric lattices.
@@ 1. Introduction
Many books have been written on the subject of matroid theory. Developed in order to generalize the concept of independent sets of vector spaces, matroids now have numerous applications in many fields. A very useful property of matroids is their large amount of cryptographic characterizations, which contributes a great deal to their many uses. Below is one such characterization using flats:
@@ 1.1. Definition:
Some of our readers may be more familiar with the following definition: a poset is a geometric lattice iff it is a finite atomic graded and submodular lattice, we encourage the readers to use [1] for more background on matroid theory .
Gelfand and Serganova [2] introduced the concept of coxeter matroids, where matroids are a special case in which the coxeter group is the symmetric group. Symplectic matroids are another important example, namely the case where the coxeter group is the hyperoctahedral group, the group of symmetries of the $n$-cube. In contrast to the case of matroids, which will be referred to from now on as ordinary matroids, we know very little about the symplectic case. In this work, we attempt to give a "lattice of flats" characterization for a class of symplectic matroids, which we call ranked. The "lattices of symplectic flats" we construct, denoted $C_{n}$ lattices, take a similar form to Definition [1.1] and have many similar properties. Our hope is that we can lay the groundwork for geometric constructions and properties in the spirit of the many simplicial complexes associated with ordinary matroids.
@@ 2. Lattice properties
We also introduce the map $*:[n]\rightarrow[n]^{*}$ defined by $i\mapsto i^{*}$ and $*:[n]^{*}\rightarrow[n]$ defined by $i^{*}\mapsto i$. Throughout this paper, we apply * to sets and collections of sets. A set $A\subset P(J)$ is called admissible if $A\cap A^{*}=\emptyset$ and a maximal admissible set will be called a transversal. We denote the collection of all admissible subsets of $J$ by $P^{ad}(J)$. The following is the main subject of this work.
@@ 2.1. Definition:
$C_{n}$ lattices are sublattices of the face lattice of the cross-polytope, resembling the ordinary case in which geometric lattices are sublattices of the face lattice of the simplex.
Before continuing with our main results, we establish some basic properties of these lattices. A partially ordered set $S$ in which every pair of elements has a unique supremum (also called join and denoted $\vee$) and a unique infimum (also called meet and denoted $\wedge$) is called a lattice. We say that a lattice is bounded if it has a greatest element (denoted $\hat{1}$) and a least element (denoted $\hat{0}$).
All $C_{n}$ lattices are bounded by $\emptyset$ and $J$. For the meet and join, we have :
@@
We say $B$ covers $A$ and denote $A\lessdot B$ if $A<B$ and for all $C$ such that $A\leq C\leq B$ we have $C=B$ or $C=A$. $A$ is an atom if it covers the least element $\hat{0}$. A bounded lattice is called atomic if every element is a join of atoms. To prove $C_{n}$ lattices are atomic we show every element $A\in\mathcal{L}$ is the join of all the atoms contained in it. We have the following:
@@
As each $A_{i}$ is contained in $A$. We now use property (4) to get that every $x\in J$ is in some atom (as they partition $J\setminus\emptyset^{*}=J$). So if $x\in A$ we have some atom $B$ s.t. $x\in B$. If $B\setminus A\neq\emptyset$ we will have $\emptyset\subsetneq B\cap A\subsetneq B$ which is a contradiction to $B$ being an atom and so $B\subseteq A$ and we have:
@@
we finish this section by showing that $C_{n}$ lattices are graded. In Section 3 we will prove that there is an underlying family of independent sets which determines this rank function. This should remind our readers of the ordinary case.
existence of rank function:
@@
Let $\mathcal{L}$ be a $C_{n}$ lattice, the existence of a rank function is equivalent to having all maximal chains between every two elements $A,B\in\mathcal{L}$ such that $A<B$ be of the same length. We prove this by induction on the maximal length of a maximal chain.
For length $1$ this is obvious as there is no chain of length 00. For some arbitrary element $A\in\mathcal{L}$ assume we have two maximal chains $\mathcal{X}$ and $\mathcal{Y}$. We can assume $\mathcal{X}\cap\mathcal{Y}=\left\{A,B\right\}$ as otherwise we are reduced to looking at maximal chains from $C\in\mathcal{X}\cap\mathcal{Y}$ to $B$ by the induction hypothesis.
From this point we will assume $A=\emptyset$, as we will prove in Section $4$ independently that the set of elements containing $A$ is order isomorphic to a $C_{n}$ lattice. We now show that every element of $\mathcal{X}$ is disjoint from every element in $\mathcal{Y}$. Assuming that this is not the case, and so they both contain an element containing some atom $C$ (since every element is a join of atoms) we can then produce chains of the same length as $\mathcal{X},\mathcal{Y}$ from $\emptyset$ to $B$ containing $C$. This will obviously be a contradiction to the induction hypothesis as we now get two chains of different length from $C$ to $B$.
@@
Hence, we can build a maximal chain from $\emptyset$ to $F$ containing $C^{\prime}$ denoted $Z_{2}$. After finitely many steps of the same process, we get a maximal chain from $\emptyset$ to $B$ going through $C^{\prime}$ with length larger than or equal to the length of $\mathcal{X}$, and since the length of $\mathcal{X}$ is strictly larger than the length of $\mathcal{Y}$ we get a contradiction.
3. Independent sets of a lattice
In [9] Andreas Blass and Bruce E. Sagan introduced the concept of $NBB$ sets as a way to calculate the möbius function of a finite bounded lattice. These sets can be thought of as a generalization of $NBC$ sets for ordinary matroids. In this section, we show that if these $NBB$ sets form a matroid we can order embed the lattice in a geometric lattice, respecting the rank function of the original lattice.
3.1. Definition:
Let $(\mathcal{L},\omega)$ be a pair of a finite bounded lattice $\mathcal{L}$ and a partial order $\omega$ on its set of atoms $A(\mathcal{L})$. A nonempty set $D\subseteq A(\mathcal{L})$ of atoms is bounded below or $BB$ if there exist $a\in A(\mathcal{L})$ such that $a$ is a strict lower bound for all $d\in D$ in the order $\omega$ and $a\leq\vee D$ in $\mathcal{L}$.
A set $B\subseteq A(\mathcal{L})$ is called $NBB$ if $B$ does not contain any bounded below subset.
Given a bounded lattice $\mathcal{L}$ and $P$ the set of all linear orders on $A(\mathcal{L})$ we define the family of independent sets $I\left(\mathcal{L},P\right)$ of $\mathcal{L}$ as follows:
$I\left(\mathcal{L},P\right)=\left\{A\in P\left(A(\mathcal{L})\right)\mid\exists\omega\in P\text{ s.t. }A\text{ is an }NBB\text{ set in }\left(\mathcal{L},\omega\right)\right\}$
We refer to elements in $I\left(\mathcal{L},P\right)$ as independent sets of $\mathcal{L}$, note that sets containing at most two atoms are always independent. The constructions in the section can be thought of as a generalization of the following observation:
@@ 3.2. Theorem:
If $\mathcal{L}$ is a geometric lattice, then $I\left(\mathcal{L},P\right)$ is the family of independent sets of the ordinary matroid corresponding to $\mathcal{L}$.
For the remainder of this section, we fix a finite, bounded, atomic, and graded lattice $\mathcal{L}$. If $\vee A=X\in\mathcal{L}$ for a set of atoms $A\subseteq A(\mathcal{L})$ we say that $A$ spans $X$ and set $rank(A)\coloneqq rank(X)$.
@@ 3.3. Lemma:
If $I$ is an independent set then there exist an atom $a\in I$ such that $rank(I\setminus\left\{a\right\})\lneq rank(I)$.
@@
If it were true that $rank(I\setminus\left\{a\right\})=rank(I)$ for every atom $a\in I$ then by our previous discussion:
$a\leq\vee\left(I\setminus\left\{a\right\}\right),\,\forall a\in I$
Let $\omega\in P$ be a linear order for which $I$ is $NBB$ and let $a$ be the first element of $I$ with respect to $\omega$. we have $a\in\vee(I\setminus\left\{a\right\})$ and so $I\setminus\left\{a\right\}$ is $BB$ by $a$ with respect to $\omega$, which is a contradiction to $I$ being an independent set.
@@ 3.4. Lemma:
The theorem follows from Lemma [3.3] and an induction argument on the rank of $I$.
In the case of geometric lattices, the rank of a flat is the size of a maximal independent set contained in it. The next two lemmas establish the same property for our lattice $\mathcal{L}$. For this statement to be true we first observe that if $I$ is an independent set and $a\in A(\mathcal{L})$ such that $rank(I)\lneq rank(I\cup\left\{a\right\})$ then $I\cup\left\{a\right\}$ is also independent, which follows again by taking the linear order for which $I$ is $NBB$ and placing $a$ first. This observation implies that every element of $\mathcal{L}$ is spanned by some independent set.
1/2/3/412341/2/3/41 / 2 / 3 / $41/23/$ 414/2/3142314/2/314 / 2 / $31/24/$ 313/2/4132413/2/413 / 2 / $412/3/$ 41/2/3412341/2/341 / 2 / 3414/23142314/2314 / 231/23412341/2341 / $3124/$ 313/24132413/2413 / 24123/41234123/4123 / 4134/21342134/2134 / 212/34123412/3412 / 341234123412341234$(a)$
[…]
3.5. Figure:
@@ 3.6. Lemma:
The rank of an element $X\in\mathcal{L}$ is the maximum size of a spanning independent set $I$.
@@
As $\vee I\leq X$, we must have $\left|I\right|\leq rank(I)\leq rank(X)$. Next we find a spanning independent set of $X$ of size $rank(X)$. Let $Y\lessdot X$, as $\mathcal{L}$ is atomic, there exists an atom $a\leq X$ such that $Y\vee a=X$. By an induction argument, we have an independent set $J$ spanning $Y$ with $\left|J\right|=rank(Y)=n$ and as $a\nleq\vee J$ we must have $I=J\cup\left\{a\right\}$ an independent set of size $n+1$ spanning $X$.
As we mentioned before, in the case of geometric lattices the rank of an element is the size of a maximal independent set contained in it. In the other direction, starting with the family of independent set, the elements of the lattice (called flats and ordered by inclusion) are the closed subsets. Formally, $F$ is a flat of the matroid $\mathcal{I}$ on the ground set $[n]$ if $rank(F)\lneq rank(F\cup\left\{i\right\})$ for all $i\notin F$.

### v6

@@ Abstract.
We are interested in expanding our understanding of symplectic matroids by exploring the properties of a class of symplectic matroids with a "lattice of flats". Taking a well-behaved family of subdivisions of the cross polytope we obtain a construction of lattices, resembling a known definition for the geometric lattice corresponding to ordinary matroid. We construct a correspondence to a set of enveloped symplectic matroids, we denote ranked symplectic matroids. As a by-product of our construction, we also obtain a new way of finding symplectic matroids from ordinary ones and an embedding Theorem into geometric lattices.
@@ 1. Introduction
Many books have been written on the subject of matroid theory. We recommend "Matroid Theory" by James Oxley for an introduction to the basics of matroid theory and "Coxeter Matroids" by Borovik, Gelfand, and White for a closer look at symplectic matroids. Developed in order to generalize the concept of independent sets of vector spaces, matroids now have numerous applications in many fields. Some of them include "Information Theory", [11], "Rigidity Theory", [12], and a recent development in algebraic geometry that helped inspire this work, [13]. A very useful property of matroids is their large amount of cryptographic characterizations, which contributes a great deal to their many uses. Below is one such characterization using flats:
@@ 1.1. Definition:
Some of our readers may be more familiar with the following definition: a poset is a geometric lattice if and only if it is a finite, atomic, graded and submodular lattice; we encourage the readers to use [1] for more background on matroid theory.
Gelfand and Serganova [2] introduced the concept of coxeter matroids, where matroids are a special case in which the coxeter group is the symmetric group. Symplectic matroids are another important example, namely the case where the coxeter group is the hyperoctahedral group, the group of symmetries of the $n$-cube. In contrast to the case of matroids, which will be referred to from now on as ordinary matroids, we know very little about the symplectic case. In this work, we attempt to give a "lattice of flats" characterization for a class of symplectic matroids, which we call ranked. The "lattices of symplectic flats" we construct, denoted $C_{n}$ lattices, take a similar form to Definition 1.1 and have many similar properties. Our hope is that we can lay the groundwork for geometric constructions and properties in the spirit of the many simplicial complexes associated with ordinary matroids.
@@ 2. Lattice properties
We also introduce the map $n]\rightarrow[n]^{*}$ defined by $i\mapsto i^{*}$ and $n]^{*}\rightarrow[n]$ defined by $i^{*}\mapsto i$. Throughout this paper, we apply * to sets and collections of sets.
2.1. Definition:
A set $A\subset P(J)$ is called admissible if $A\cap A^{*}=\emptyset$ and a maximal admissible set will be called a transversal, in other words, a transversal contains precisely one copy of $i,i^{*}$ for each $i\in[n]$. We denote the collection of all admissible subsets of $J$ by $P^{ad}(J)$.
2.2. Definition:
For a lattice $\mathcal{L}$ and $A,B\in\mathcal{L}$ we say that $B$ covers $A$ and denote $A\lessdot B$ if $A<B$ and for all $C\in\mathcal{L}$ such that $A\leq C\leq B$ we have $C=B$ or $C=A$. $A\in\mathcal{L}$ is an atom if it covers the least element $\hat{0}$.
The following is the main subject of this work.
@@ 2.3. Definition:
$C_{n}$ lattices are subsets of the face lattice of the cross-polytope, resembling the ordinary case in which geometric lattices are subsets of the face lattice of the simplex. This gives rise to a neutral realization of $C_{n}$ lattices.
Example:
The simplest example of a $C_{n}$ lattice is $\mathcal{L}(U^{*}_{k,n})$, the collection of admissible sets of size no greater than k1k-1.
Before continuing with our main results, we establish some basic properties of $C_{n}$ lattices. A partially ordered set $S$ in which every pair of elements has a unique supremum (also called join and denoted $\vee$) and a unique infimum (also called meet and denoted $\wedge$) is called a lattice. Every finite lattice contains a greatest element (denoted $\hat{1}$) and a least element (denoted $\hat{0}$).
For a $C_{n}$ lattice we have the following meet and join :
@@
As $C_{n}$ lattices are a finite poset with a meet, they are indeed lattices.
2.4. Lemma:
The atoms in a $C_{n}$ lattice partition $J$.
Proof:
Using property (1) in 2.3, the least element of a geometric lattice is always the empty set. The Lemma now follows directly from property (4) in Definition 2.3.
2.5. Definition:
A finite lattice is called atomic if every element is a join of atoms.
2.6. Theorem:
$C_{n}$ lattices are atomic.
Proof:
Let $\mathcal{L}$ be a $C_{n}$ lattice, we will show that every element $A\in\mathcal{L}$ is the join of all the atoms contained in it. We have the following:
@@ 2.6. Theorem:
As each $A_{i}$ is contained in $A$. We now use Lemma 2.4 to see that every $x\in J$ is in some atom. So, if $x\in A$ we have some atom $B$ s.t. $x\in B$. If $B\setminus A\neq\emptyset$ we will have $\emptyset\subsetneq B\cap A\subsetneq B$ which is a contradiction to $B$ being an atom and so $B\subseteq A$ and we have:
@@
We will now show that $C_{n}$ lattices are graded. In Section $3$ we will prove that there is an underlying family of independent sets which determines this rank function. This should remind our readers of the ordinary case.
2.7. Definition:
@@
A lattice $\mathcal{L}$ is called graded if for every $A,B\in\mathcal{L}$ such that $A<B$ and for every two maximal chains $\mathcal{X}=\left\{A=X_{1}\lessdot...\lessdot X_{m}=B\right\},\mathcal{Y}=\left\{A=Y_{1}\lessdot...\lessdot Y_{l}=B\right\}$ we have $m=l$. This enables us to define the rank function $r(A)$ of a graded lattice to be the length of a maximal chain from $\hat{0}$ to $A$.
$BD\neq B$$^{\prime\prime}DD^{\prime}$$C^{\prime\prime}CC^{\prime}$AA
2.8. Figure:
The elements described in the proof of Theorem 2.9. The straight lines represent covering relations, and dashed lines represent containment.
2.9. Theorem:
$C_{n}$ lattices are graded.
Proof:
Let $\mathcal{L}$ be a $C_{n}$ lattice and $A,B\in\mathcal{L}$ such that $A<B$ . We prove the Theorem by induction on $n$ the maximum length of a maximal chain between $A$ and $B$.
For length $n=3$, this is obvious as only the empty chain is of length 0. Let $\mathcal{X},\mathcal{Y}\subseteq\mathcal{L}$ be two maximal chains between $A$ and $B$.
Let $C\lessdot D$ and $C^{\prime}\lessdot D^{\prime}$ be the first two elements other than $A$ in $\mathcal{X},\mathcal{Y}$, respectively, as illustrated in Figure 2.8. We first show that $D,D^{\prime}\neq B$ or $D=D^{\prime}=B$, the second case is not interesting as then both chains are of length $3$. Let $D=B$ and assume that there exists $a\in C\setminus C^{\prime*}$. By property (4) of Definition 2.3 we have an element $D^{\prime\prime}\in\mathcal{L}$, covering $C^{\prime}$ and containing $a$. As $a\in D^{\prime\prime}\wedge C$ and $C\lessdot D$, we have
$A\subsetneq D^{\prime\prime}\wedge C\subset C\Rightarrow C\subset D^{\prime\prime}\Rightarrow D=D^{\prime}=D^{\prime\prime}=B$
If there is no such $a$ then let $b\in D^{\prime}\setminus C^{\prime}$ If $D^{\prime}\neq B$ then $b\notin C\cup C^{*}$. Using property (4) of Definition 2.3, we obtain an element $C^{\prime\prime}\in\mathcal{L}$ covering $A$ and containing $b$. We have now reduced the problem to the first case, replacing $C^{\prime}$ with $C^{\prime\prime}$.
To finish the proof of the Theorem, We now first observe the case $(C\setminus A)^{*}\subset C^{\prime}$, illustrated in Figure 2.8. If that is indeed the case, let $a\in D^{\prime}\setminus C^{\prime}$ and define $C^{\prime\prime}$ as the element that covers $A$ and contains $a$ resulting from property (4) of Definition 2.3. As $\mathcal{L}$ is atomic we have $C^{\prime\prime}<D^{\prime}$. By applying the induction hypotheses, using the fact $D^{\prime}\neq B$, to the chains $\left\{A,C^{\prime},D^{\prime}\right\}$ and $\left\{A,C^{\prime\prime},D^{\prime}\right\}$, we actually have $C^{\prime\prime}\lessdot D^{\prime}$. Once again using property (4) of Definition 2.3, we have $C^{\prime}\cap C^{\prime\prime}=A$, therefore for each $b\in C\setminus A$ there is an element $D^{\prime\prime}\in\mathcal{L}$ covering $C^{\prime\prime}$ and containing $b$. We trivially have
$A<D^{\prime\prime}<B\Rightarrow A\lneq C\wedge D^{\prime\prime}<C\Rightarrow C<D^{\prime\prime}$
We again apply the induction hypotheses, using the fact that $D^{\prime\prime}\neq B$, to the chains $\left\{A,C,D^{\prime\prime}\right\}$ and $\left\{A,C^{\prime\prime},D^{\prime\prime}\right\}$ to obtain $C\lessdot D^{\prime\prime}$.
Let $\mathcal{Z}\subseteq\mathcal{L}$ be some maximal chain from $A$ to $B$ with $C^{\prime\prime}\lessdot D^{\prime\prime}$, its first two elements other than $A$.
To show $\mathcal{X}$ is of the same length as $\mathcal{Z}$ we apply the induction hypothesis to $\mathcal{X}\setminus\left\{A\right\}$ and $(\mathcal{Z}\setminus\left\{A,C^{\prime\prime}\right\})\cup\left\{A,C\right\}$. The same is done to show $\mathcal{Y}$ is of the same length as $\mathcal{Z}$, producing the proof. Finally, We observe that if $(C\setminus A)^{*}\setminus C^{\prime}\neq\emptyset$ we can choose $C^{\prime\prime}=c^{\prime}$ and $D^{\prime\prime}$ to be an element covering $C^{\prime}$ and containing some element in $(C\setminus A)^{*}\setminus C^{\prime}$. The rest of the proof is similar.
To finish this section we show that $C_{n}$ lattices are "almost" geometric, with only one missing property, submodularity.
2.10. Definition:
A graded lattice $\mathcal{L}$ will be called submodular if for every $A,B\in\mathcal{L}$ its rank function obeys the identity:
$rank(A)+rank(B)\geq rank(A\vee B)+rank(A\wedge B)$
2.11. Theorem:
A $C_{n}$ lattice $\mathcal{L}$ is a geometric lattice if and only if $rank(\mathcal{L})\leq 2$
Proof:
If $rank(\mathcal{L})\leq 2$, then property (4) of Definition 2.3 implies property (3) of Definition 1.1. As the other properties of Definition 1.1 are trivially true, we have $\mathcal{L}$ a geometric lattice. If $rank(\mathcal{L})\geq 2$, then there always exist two atoms $A,B$ of $\mathcal{L}$ such that $1\in A$ and $1^{*}\in B$. We have $A\vee B=J,A\wedge B=\emptyset$ and so:
$rank(A)+rank(B)=2\lneq rank(\mathcal{L})=rank(J)+rank(\emptyset)=rank(A\vee B)+rank(A\wedge B)$
Therefore, $\mathcal{L}$ is not submodular and is not a geometric lattice.
3. Independent sets of a lattice
In [9] Andreas Blass and Bruce E. Sagan introduced the concept of $NBB$ sets as a way to calculate the möbius function of a finite lattice. These sets can be thought of as a generalization of $NBC$ sets for ordinary matroids. In this section, we show that if these $NBB$ sets form a matroid, we can order embed the lattice in a geometric lattice, respecting the rank function of the original lattice. We make a small change in the definition of $BB$ sets found in [9] by considering only linear orders on the atom set of the lattice, resulting in a simpler definition.
3.1. Definition:
Let $(\mathcal{L},\omega)$ be a pair of a finite lattice $\mathcal{L}$ and a linear order $\omega$ on its set of atoms $A(\mathcal{L})$. A nonempty set $D\subseteq A(\mathcal{L})$ of atoms is bounded below or $BB$ if there exist $a\in A(\mathcal{L})$ such that $a$ precedes $D$ in the order $\omega$ and $a\leq\vee D$ in $\mathcal{L}$. A set $B\subseteq A(\mathcal{L})$ is called $NBB$ if $B$ does not contain any bounded below subset.
Given a finite lattice $\mathcal{L}$ and $T$ the set of all linear orders on $A(\mathcal{L})$ we define the family of independent sets $I\left(\mathcal{L}\right)$ of $\mathcal{L}$ as follows:
@@ 3.1. Definition:
We refer to the elements in $I\left(\mathcal{L}\right)$ as independent sets of $\mathcal{L}$. Before continuing, we note the following immediate properties:
  1. (1)
    A subset of an independent set is itself independent. In other words, independent sets form an abstract simplicial complex.
  2. (2)
    A set containing a single atom is not $BB$ in any order.
  3. (3)
    sets containing at most two atoms are always independent. This is done by choosing any order where one of the two elements is first.
An immediate implication of the proof of Theorem [1.2] in [9] is the following Theorem.
@@ 3.2. Theorem:
If $\mathcal{L}$ is a geometric lattice, then $I\left(\mathcal{L}\right)$ is the family of independent sets of the ordinary matroid corresponding to $\mathcal{L}$.
For the remainder of this section, we fix a finite, atomic, and graded lattice $\mathcal{L}$. If $\vee A=X\in\mathcal{L}$ for a set of atoms $A\subseteq A(\mathcal{L})$ we say that $A$ spans $X$ and define $rank(A)\coloneqq rank(X)$.
@@ 3.3. Lemma:
If $I$ is an independent set, then there exists an atom $a\in I$ such that $rank(I\setminus\left\{a\right\})\lneq rank(I)$.
@@
For any atom $a\in I$ if $rank(I\setminus\left\{a\right\})=rank(I)$ then $a\leq\vee I=\vee\left(I\setminus\left\{a\right\}\right)$.
Let $\omega$ be a linear order on $A(\mathcal{L})$ for which $I$ is $NBB$ and let $a$ be the first element of $I$ with respect to $\omega$. we have $a\in\vee(I\setminus\left\{a\right\})$ and so $I\setminus\left\{a\right\}$ is $BB$ by $a$ with respect to $\omega$, which is a contradiction to $I$ being $NBB$ for $\omega$.
@@ 3.4. Lemma:
The Theorem follows from Lemma 3.3 and an induction argument on the rank of $I$.
In the case of geometric lattices, the rank of a flat is the size of a maximal independent set contained in it. The next two lemmas establish the same property for our lattice $\mathcal{L}$.
@@ 3.5. Lemma:
Let $I$ be an independent set and $a\in A(\mathcal{L})$ such that $rank(I)\lneq rank(I\cup\left\{a\right\})$ then $I\cup\left\{a\right\}$ is also independent.
@@
Let $\omega$ be a linear order on $A(\mathcal{L})$ for which $I$ is $NBB$. Define $\omega_{a}$ as the linear order obtained from $\omega$ by placing $a$ first. checking the conditions of Definition 3.1 we see that $I$ is $NBB$ with respect to $\omega_{a}$ as $a\nleq\vee I$. Therefore, $I\cup\left\{a\right\}$ is $NBB$ with respect to $\omega_{a}$ as adding the first element does not create new $BB$ subsets.
3.6. Lemma:
The rank of an element $X\in\mathcal{L}$ is the cardinality of a maximal independent set $I$ of atoms smaller than $X$.
Proof:
As all atoms in $I$ are smaller than $X$, we have $\vee I\leq X$, hence Lemma 3.4 implies $|I|\leq rank(I)\leq rank(X)$. Next, we find an independent set of $X$ of size $rank(X)$. Let $Y\lessdot X$, as $\mathcal{L}$ is atomic, there exists an atom $a\leq X$ such that $Y\vee a=X$. By an induction argument, we have an independent set $J$ contained in $Y$ with $\left|J\right|=rank(Y)=rank(X)-1$ and as $a\nleq\vee J$ we must have $I=J\cup\left\{a\right\}$ an independent set of size $rank(X)$ contained in $X$.
As noted previously, in geometric lattices, the rank of an element is the cardinality of a maximal independent set that it contains. Beginning with the collection of independent sets of a matroid, the elements of its associated geometric lattice - termed flats and ordered by inclusion - are precisely its closed subsets. More formally, a subset $F$ is a flat of the matroid $\mathcal{M}$ (i.e., an element of its geometric lattice) if and only if $rank(F)\lneq rank(F\cup\left\{e\right\})$ for every element $e\notin F$.